OpenStudy (anonymous):

I'm still new to these types of mathematics, I need help... One car traveled 248 miles and drove 5 mph hour slower than a second car which drove 268 miles. If the cars were traveling for the same time, how fast was each car traveling? Show all work. Show all work.

OpenStudy (anonymous):

x = rate rt = d or xt=d d/x = t 248/x-5 = 268/x

OpenStudy (anonymous):

248x = 268(x-5) Solve for x.

OpenStudy (anonymous):

248x = 268x - 1340 248x - 268x = -1340 -20x = -1340 x = 67 One car was going 67, the other was going 62

OpenStudy (anonymous):

Thank you SO much!